OpenStudy (chemicaldiscoboy):

write the polynomial in standard form -9x + x^3 - x^4 + 3

OpenStudy (triciaal):

standard form: the variable in descending order

OpenStudy (chemicaldiscoboy):

So do I reorder it in descending order?

OpenStudy (3mar):

we have four terms: -9x x^3 -x^4 3 and the polynomial in standard form looks like that: \[\left[ ? \right]x^4+\left| ? \right|x^3+\left| ? \right|x^2+\left| ? \right|x^1+\left| ? \right|x^0\]

OpenStudy (chemicaldiscoboy):

So if I write it in descending order it would be -x^4 + x^3 - 9x + 3?

OpenStudy (3mar):

Great!

OpenStudy (chemicaldiscoboy):

Thanks for the help!

OpenStudy (3mar):

You are welcome! Any Help... Anytime.

OpenStudy (3mar):

Any more questions?

OpenStudy (3mar):

Thank you for the medal!